# Bayesian Ego-graph inference for Networked Multi-Agent Reinforcement Learning

### v1

@@ Abstract
In networked multi-agent reinforcement learning (Networked-MARL), decentralized agents must act under local observability and constrained communication over fixed physical graphs. Existing methods often assume static neighborhoods, limiting adaptability to dynamic or heterogeneous environments. While centralized frameworks can learn dynamic graphs, their reliance on global state access and centralized infrastructure is impractical in real-world decentralized systems. We propose a stochastic graph-based policy for Networked-MARL, where each agent conditions its decision on a sampled subgraph over its local physical neighborhood. Building on this formulation, we introduce BayesG, a decentralized actor–critic framework that learns sparse, context-aware interaction structures via Bayesian variational inference. Each agent operates over an ego-graph and samples a latent communication mask to guide message passing and policy computation. The variational distribution is trained end-to-end alongside the policy using an evidence lower bound (ELBO) objective, enabling agents to jointly learn both interaction topology and decision-making strategies. BayesG outperforms strong MARL baselines on large-scale traffic control tasks with up to 167 agents, demonstrating superior scalability, efficiency, and performance.
@@ 2 Related work
Graph-based MARL and Latent Structure Learning. Recent advances in cooperative MARL leverage latent interaction graphs to enhance coordination and scalability. Methods such as DGN [22], DICG [23], and HGAP [24] employ attention mechanisms or graph neural networks (GNNs) [25, 26] to infer inter-agent dependencies and guide information exchange. Variational approaches [19, 21] further treat interaction graphs as latent variables, jointly optimizing communication structure and policy learning. However, most existing methods assume centralized training, access to global state, and unrestricted graph rewiring [13, 14, 20], which are impractical in real-world decentralized settings. In contrast, we address the more realistic networked MARL scenario by learning dynamic latent interaction masks on the environment graph using only local observations.
Networked MARL. Networked MARL focuses on decentralized learning over fixed topologies, often without explicitly adapting the communication structure. Consensus-based methods [15, 16] synchronize local value functions via neighborhood averaging, typically assuming partial or global observability. Other works incorporate network-aware priors such as distance-based decay [27] or local reward aggregation [18] to promote spatially coherent coordination. Model-based frameworks [28, 29] exchange predicted trajectories but still rely on static interaction graphs. Communication-based methods like NeurComm [7] enable information sharing across neighbors, yet use fixed topologies throughout training. Our method differs fundamentally by introducing a Bayesian latent graph inference mechanism that allows each agent to adaptively select task-relevant neighbors from its physical graph, enabling communication-efficient, locally grounded coordination.
@@ 5.1 Experimental Setup
We evaluate BayesG on five benchmark scenarios for adaptive traffic signal control (ATSC), implemented using the SUMO microscopic traffic simulator [32]. Each environment simulates peak-hour traffic, with one MDP step corresponding to a fixed control interval. Agents observe real-time traffic conditions via induction-loop detectors (ILDs), including vehicle density, queue lengths, and waiting times on incoming lanes. Each agent controls a local traffic signal and receives both local and neighbor-level statistics. The reward is defined as the negative number of halted vehicles on incoming lanes, normalized by a fixed scale—encouraging queue minimization.
@@
NewYork. To assess scalability, we introduce three large-scale scenarios—NewYork33, NewYork51, and NewYork167—comprising 33, 51, and 167 signalized intersections, respectively. These networks are derived from real-world Manhattan layouts (See Appendix B for more details). We use a 20-second control interval (increased to 40 seconds for NewYork167) and simulate 500 steps per episode. Intersections feature heterogeneous phase designs and ILD configurations. States include normalized lane-level metrics and neighborhood-aware features. These environments pose significant challenges for coordination and generalization in decentralized MARL.
@@ 5.1.2 Implementation Details
$\tilde{s}_{i}=\mathrm{GNN}_{\text{obs}}(S_{\mathcal{V}_{i}},A_{i}^{*})\,\|\,\mathrm{GNN}_{\text{policy}}(\Pi_{\mathcal{V}_{i}},A_{i}^{*})\,\|\,\mathrm{GNN}_{\text{traj}}(H_{\mathcal{V}_{i}},A_{i}^{*})$ (12)
Here, $S_{\mathcal{V}_{i}}$, $\Pi_{\mathcal{V}_{i}}$, and $H_{\mathcal{V}_{i}}$ represent the observation, policy, and trajectory features of agent $i$’s neighborhood. While any GNN can be employed, we use graph convolutional networks (GCNs) [36] for efficiency in our implementation. This formulation allows each agent to learn sparse and context-aware communication structures tailored to local dynamics. (see pseudocode in Appendix C).
@@ 5.2 Performance Comparison
To further understand the optimization dynamics of BayesG, we report the evolution of key training loss components in Appendix D for ATSC_Grid, Monaco, and NewYork33. Specifically, we track the policy loss, value (critic) loss, ELBO loss, prior loss, mask entropy regularization, and the total objective. These loss curves offer deeper insight into how BayesG balances policy optimization with variational graph inference, and demonstrate the stability of the joint training process across different network scales.
Figure 3: Visualization of traffic congestion on the Grid map at 3500 simulation seconds. Road thickness and color represent vehicle density (thicker and redder indicates more congestion). BayesG results in significantly lighter and sparser traffic flows.Additional qualitative comparisons across different simulation times are provided in Figure 10 (Appendix E).
Qualitative Comparison on Grid Environment. Figure 3 presents a visual comparison of traffic congestion for IA2C, CommNet, NeurComm, and BayesG on the ATSC_Grid map at 3,500 simulation seconds. Road segment thickness and color indicate vehicle density, with thicker and redder lines denoting heavier congestion. IA2C and CommNet exhibit severe congestion across many intersections, while NeurComm reduces some bottlenecks through explicit communication. BayesG, in contrast, achieves the smoothest traffic flow, with thinner and less congested segments. These results highlight BayesG’s ability to adaptively focus on critical interactions via its learned latent graph, enabling more effective and scalable coordination in dense traffic scenarios. Additional qualitative comparisons across different simulation times are provided in Appendix E.
@@ 5.3 Case Study: Interpreting Learned Interaction Structures
This visualization shows that BayesG learns to prioritize communication from low-congestion intersections toward more congested regions. For example, when congestion builds up in a particular area (e.g., top-right and bottom right clusters), left-stream intersections increase their coordination weights, effectively regulating inflow and providing more green phases to help alleviate downstream pressure. Unlike approaches that treat all neighbors equally, BayesG captures directional, task-driven cooperation: it encourages agents outside a congested zone to adjust proactively, thereby relieving bottlenecks before they worsen. This reflects real-world traffic control principles, where strategic upstream coordination plays a critical role in minimizing congestion spread. These results demonstrate BayesG’s ability to infer decentralized, context-aware communication structures that adapt dynamically to evolving traffic conditions. (See more cases in Appendix F)
@@ Acknowledgements
This work is supported by the Australian Research Council under Australian Laureate Fellowships FL190100149 and Discovery Early Career Researcher Award DE200100245.
@@ 3. Entropy Term: $\log q(Z_{i};\phi_{i})$
Appendix B NewYork Scenario Visualization and Graph Statistics
@@ Appendix C Algorithms
Appendix D Training Loss Analysis
@@
Appendix E Additional Visualizations on Grid Environment
@@
Appendix F Extended Case Study: Dynamic Graph Adaptation Across Time and Runs
@@
Appendix G Limitations
@@
  • Evaluation domain. Our experiments focus on adaptive traffic signal control, which is a structured and spatially grounded application. Although the method is broadly applicable to networked MARL, its performance on less structured or non-spatial tasks (e.g., multi-robot coordination, communication in games) has not been explored [40, 41].

### v2

@@ Abstract
In networked multi-agent reinforcement learning (Networked-MARL), decentralized agents must act autonomously under local observability and constrained communication over fixed physical graphs. Existing methods often assume static neighborhoods, limiting adaptability to dynamic or heterogeneous environments. While centralized frameworks can learn dynamic graphs, their reliance on global state access and centralized infrastructure is impractical in real-world decentralized systems. We propose a stochastic graph-based policy for Networked-MARL, where each agent conditions its decision on a sampled subgraph over its local physical neighborhood. Building on this formulation, we introduce BayesG, a decentralized actor–critic framework that learns sparse, context-aware interaction structures via Bayesian variational inference. Each agent operates over an ego-graph and samples a latent communication mask to guide message passing and policy computation. The variational distribution is trained end-to-end alongside the policy using an evidence lower bound (ELBO) objective, enabling agents to jointly learn both interaction topology and decision-making strategies. BayesG outperforms strong MARL baselines on large-scale traffic control tasks with up to 167 agents, demonstrating superior scalability, efficiency, and performance.
@@ 2 Related work
Networked MARL. Networked MARL focuses on decentralized learning over fixed topologies, often without explicitly adapting the communication structure. Consensus-based methods [15, 16] synchronize local value functions via neighborhood averaging, typically assuming partial or global observability. Other works incorporate network-aware priors such as distance-based decay [22] or local reward aggregation [18] to promote spatially coherent coordination. Model-based frameworks [23, 24] exchange predicted trajectories but still rely on static interaction graphs. Communication-based methods like NeurComm [7] enable information sharing across neighbors, yet use fixed topologies throughout training. Recent work [25, 22, 26] has explored sampling strategies to improve scalability in networked systems, but these methods sample from predefined or uniform distributions at the environment or agent level. Our method differs fundamentally by introducing a Bayesian latent graph inference mechanism that allows each agent to adaptively select task-relevant neighbors from its physical graph, enabling communication-efficient, locally grounded coordination.
Graph-based MARL and Latent Structure Learning. Recent advances in cooperative MARL leverage latent interaction graphs to enhance coordination and scalability. Methods such as DGN [27], DICG [28], and HGAP [29] employ attention mechanisms or graph neural networks (GNNs) [30, 31] to infer inter-agent dependencies and guide information exchange. Variational approaches [19, 21] further treat interaction graphs as latent variables, jointly optimizing communication structure and policy learning. However, most existing methods, including Dec-POMDP-based coordination graph approaches [32, 28, 33, 13, 20], assume centralized training with access to global state and reward signals, and allow unrestricted graph rewiring, which are impractical in real-world decentralized settings. In contrast, we address the more realistic networked MARL scenario by learning dynamic latent interaction masks on the environment graph using only local observations.
@@ 4.2 BayesG: Variational Latent Graph Inference for Policy Optimisation
Treating edge masks as latent variables enables agents to capture uncertainty over which neighbors are informative—particularly valuable when noisy or non-stationary dynamics make deterministic graph pruning unreliable.
@@
Connection to maximum entropy RL. Our formulation parallels the probabilistic interpretation of RL in Soft Actor-Critic (SAC) [36], which casts policy optimization as entropy-regularized inference. In SAC, policy entropy $\mathcal{H}(\pi(\cdot|s))$ regularizes action selection; in our framework, the term $\log q(Z_{i};\phi_{i})$ in Equation (8) acts as mask entropy $\mathcal{H}(q(Z_{i};\phi_{i}))$, regularizing the variational distribution over communication structures. This prevents premature collapse to deterministic graphs and enables uncertainty-aware neighbor selection. The dual entropy regularization—over both actions (embedded in $\mathcal{L}_{\theta,\varphi}$) and edges (the $\log q$ term)—is key to our method’s robustness in dynamic environments where action diversity and adaptive communication are both critical.
@@ 5.1 Experimental Setup
We evaluate BayesG on five benchmark scenarios for adaptive traffic signal control (ATSC), implemented using the SUMO microscopic traffic simulator [37]. Each environment simulates peak-hour traffic, with one MDP step corresponding to a fixed control interval. Agents observe traffic conditions via induction-loop detectors (ILDs), including vehicle density, queue lengths, and waiting times on incoming lanes, and control local traffic signals. The reward is the negative number of halted vehicles, normalized by a fixed scale. These environments naturally satisfy Definition 1: state transitions depend on local and neighbor actions, rewards are localized per agent, and neighborhood structure is fixed by the road network. Detailed MDP mappings are in Appendix B.
@@
NewYork. To assess scalability, we introduce three large-scale scenarios—NewYork33, NewYork51, and NewYork167—comprising 33, 51, and 167 signalized intersections, respectively. These networks are derived from real-world Manhattan layouts (See Appendix C for more details). We use a 20-second control interval (increased to 40 seconds for NewYork167) and simulate 500 steps per episode. Intersections feature heterogeneous phase designs and ILD configurations. States include normalized lane-level metrics and neighborhood-aware features. These environments pose significant challenges for coordination and generalization in decentralized MARL.
@@ 5.1.2 Implementation Details
$\tilde{s}_{i}=\mathrm{GNN}_{\text{obs}}(S_{\mathcal{V}_{i}},A_{i}^{*})\,\|\,\mathrm{GNN}_{\text{policy}}(\Pi_{\mathcal{V}_{i}},A_{i}^{*})\,\|\,\mathrm{GNN}_{\text{traj}}(H_{\mathcal{V}_{i}},A_{i}^{*}).$ (12)
Here, $S_{\mathcal{V}_{i}}$, $\Pi_{\mathcal{V}_{i}}$, and $H_{\mathcal{V}_{i}}$ represent the observation, policy, and trajectory features of agent $i$’s neighborhood. While any GNN can be employed, we use graph convolutional networks (GCNs) [41] for efficiency in our implementation. This formulation allows each agent to learn sparse and context-aware communication structures tailored to local dynamics. (see pseudocode in Appendix D).
@@ 5.2 Performance Comparison
To further understand the optimization dynamics of BayesG, we report the evolution of key training loss components in Appendix E for ATSC_Grid, Monaco, and NewYork33. Specifically, we track the policy loss, value (critic) loss, ELBO loss, prior loss, mask entropy regularization, and the total objective. These loss curves offer deeper insight into how BayesG balances policy optimization with variational graph inference, and demonstrate the stability of the joint training process across different network scales.
Figure 3: Visualization of traffic congestion on the Grid map at 3500 simulation seconds. Road thickness and color represent vehicle density (thicker and redder indicates more congestion). BayesG results in significantly lighter and sparser traffic flows.Additional qualitative comparisons across different simulation times are provided in Figure 10 (Appendix F).
Qualitative Comparison on Grid Environment. Figure 3 presents a visual comparison of traffic congestion for IA2C, CommNet, NeurComm, and BayesG on the ATSC_Grid map at 3,500 simulation seconds. Road segment thickness and color indicate vehicle density, with thicker and redder lines denoting heavier congestion. IA2C and CommNet exhibit severe congestion across many intersections, while NeurComm reduces some bottlenecks through explicit communication. BayesG, in contrast, achieves the smoothest traffic flow, with thinner and less congested segments. These results highlight BayesG’s ability to adaptively focus on critical interactions via its learned latent graph, enabling more effective and scalable coordination in dense traffic scenarios. Additional qualitative comparisons across different simulation times are provided in Appendix F.
@@ 5.3 Case Study: Interpreting Learned Interaction Structures
This visualization shows that BayesG learns to prioritize communication from low-congestion intersections toward more congested regions. For example, when congestion builds up in a particular area (e.g., top-right and bottom right clusters), left-stream intersections increase their coordination weights, effectively regulating inflow and providing more green phases to help alleviate downstream pressure. Unlike approaches that treat all neighbors equally, BayesG captures directional, task-driven cooperation: it encourages agents outside a congested zone to adjust proactively, thereby relieving bottlenecks before they worsen. This reflects real-world traffic control principles, where strategic upstream coordination plays a critical role in minimizing congestion spread. These results demonstrate BayesG’s ability to infer decentralized, context-aware communication structures that adapt dynamically to evolving traffic conditions. (See more cases in Appendix G)
@@ 5.4 Ablation Studies
6 Discussion
6.1 Comparison with Sampling-based Networked MARL
Recent work has explored sampling strategies for scalability in networked MARL, but differs fundamentally in what is sampled, how distributions are determined, and where adaptivity occurs.
Sampling targets and distributions. Qu et al. [25] establish convergence results for networked systems with local dependencies on fixed graphs. Lin et al. [22] sample active links from a predefined distribution $\mathcal{D}$ for $\mu$-decay analysis. Anand and Qu [26] use uniform agent subsampling ($k$ out of $n$) for a global controller with mean-field bounds. In contrast, our method learns context-dependent, per-edge distributions: each agent maintains a learned posterior $q(Z_{i};\phi_{i})$ over its 1-hop edges, optimized jointly with the policy via an ELBO.
Locus of adaptivity. Lin et al. [22] achieve dynamics-level stochasticity through time-varying link sets with multi-hop influence across time. Anand and Qu [26] vary agent subsets for a global controller. Our approach operates at the policy/representation level: each agent samples a task-adaptive 1-hop subgraph for local decision-making.
First-hop inclusion probabilities. Lin et al. [22] use the marginal of $\mathcal{D}$ (predefined, not state-adaptive). Anand and Qu [26] use uniform inclusion ($k/n$) with no edge-specific notion. Our method learns edge-specific probabilities $\Pr[z_{ij}=1]=\sigma(\phi_{ij}(\cdot))$ that are state and trajectory-dependent.
These approaches provide valuable convergence guarantees under structural assumptions. Our contribution is complementary: learning which neighbors to attend to, optimized end-to-end with the policy. Extending their theoretical tools to learned edge masks is a promising future direction. A detailed side-by-side comparison is provided in Table 2 (Appendix H).
6.2 Distinction from Dec-POMDP-based Coordination Graphs
Our method addresses the Spatiotemporal-MDP setting (Definition 1), where agent transitions $p_{i}(s_{i}^{\prime}|s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}})$ and rewards $R(s_{\mathcal{V}_{i}},u_{\mathcal{V}_{i}})$ are localized over physically connected neighbors determined by $G_{\text{env}}$. This reflects real-world systems (traffic networks, power grids, sensor fields) where agents only observe and affect local neighborhoods through fully decentralized learning.
In contrast, Dec-POMDP-based methods such as CASEC [14], DCG [32], DICG [28], SOP-CG [33], and GACG [21] assume global state $s$, global reward $R(s,\mathbf{u})$, joint transitions $p(s^{\prime}|s,\mathbf{u})$, and centralized training with decentralized execution (CTDE). They allow unrestricted graph rewiring unconstrained by physical topology, making them incompatible with physically grounded networked systems.
Our baselines—CommNet [39], NeurComm [7], and LToS [22]—communicate only with physical neighbors, respect fixed topology, and perform fully decentralized learning. While Dec-POMDP coordination graphs are important contributions, we exclude them due to incompatibility with Spatiotemporal-MDP’s decentralized, physically constrained nature.
@@ Acknowledgements
This work is supported by the Australian Research Council under Australian Laureate Fellowships FL190100149.
@@ 3. Entropy Term: $\log q(Z_{i};\phi_{i})$
Appendix B ATSC Environment: MDP Component Mapping and Spatiotemporal-MDP Justification
We provide detailed mappings from the adaptive traffic signal control (ATSC) domain to the Spatiotemporal-MDP framework (Definition 1).
B.1 MDP Components
States ($s_{i}$). Each agent $i$ (a signalized intersection) observes its local traffic state via induction loop detectors (ILDs) placed on incoming lanes. The state includes:
  • Vehicle density: Number of vehicles per lane
  • Queue length: Number of stopped vehicles per lane
  • Waiting time: Average waiting time of vehicles per lane
  • Current phase: Active traffic signal phase (e.g., north-south green)
  • Phase duration: Time elapsed in current phase
For neighborhood-aware coordination, agents also receive aggregated statistics from immediate neighbors $\mathcal{N}_{i}$ (e.g., neighbor queue lengths, active phases).
Actions ($u_{i}$). Each agent selects a traffic signal phase from a discrete action space. Actions correspond to:
  • Phase switching: Transition to a different signal phase (e.g., from north-south green to east-west green)
  • Phase holding: Maintain the current phase for another control interval
A mandatory yellow phase (2 seconds) is enforced between conflicting green phases for safety.
Transition Probabilities ($p_{i}(s_{i}^{\prime}|s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}})$). Transitions are not analytically available but are governed by the SUMO microscopic traffic simulator [37]. The key property is locality: agent $i$’s next state $s_{i}^{\prime}$ depends primarily on:
  • Its own action $u_{i}$ (phase decision)
  • Local state $s_{i}$ (current congestion)
  • Neighbor actions $u_{\mathcal{N}_{i}}$ (upstream/downstream traffic release)
This satisfies the Spatiotemporal-MDP assumption that transitions are localized over the physical neighborhood $\mathcal{V}_{i}$.
Rewards ($R(s_{\mathcal{V}_{i}},u_{\mathcal{V}_{i}})$). The reward for agent $i$ is:
$r_{i}=-\frac{1}{C}\sum_{\ell\in\text{lanes}_{i}}\text{num\_halted}(\ell),$ (34)
where $\text{num\_halted}(\ell)$ is the number of stopped vehicles on lane $\ell$, and $C$ is a normalization constant. This encourages local queue minimization. The reward is localized to agent $i$’s intersection, consistent with the Spatiotemporal-MDP framework.
B.2 Why Spatiotemporal-MDP Fits ATSC
The ATSC domain is a canonical example of Spatiotemporal-MDP due to:
1. Localized dynamics. Traffic flow is governed by physical proximity: upstream intersections release vehicles that propagate to downstream intersections. Each agent’s state evolution depends on its immediate neighbors’ actions, not the global joint action of all agents.
2. Fixed physical topology. The road network structure is fixed and sparse, with agents (intersections) only interacting with directly connected neighbors via shared road segments.
3. Decentralized execution requirement. In real-world deployments, traffic signals operate independently with limited communication bandwidth. Centralized control is impractical due to:
  • Scalability: City-scale networks have hundreds of intersections; centralized joint action spaces grow exponentially
  • Communication constraints: Real-time global state aggregation is infeasible under latency and bandwidth limits
  • Robustness: Centralized systems are vulnerable to single points of failure
4. Local observability. Each intersection has sensors only for its incoming lanes, consistent with the partial observability assumption in Spatiotemporal-MDP.
These properties make ATSC fundamentally different from cooperative benchmarks (e.g., StarCraft) that assume global rewards, unrestricted communication, and arbitrary coordination graphs. Our method exploits this structure by learning sparse, physically grounded communication masks, enabling scalable and deployable traffic control.
Appendix C NewYork Scenario Visualization and Graph Statistics
@@ Appendix D Algorithms
Appendix E Training Loss Analysis
@@
Appendix F Additional Visualizations on Grid Environment
@@
Appendix G Extended Case Study: Dynamic Graph Adaptation Across Time and Runs
@@
Appendix H Detailed Comparison with Sampling-based Networked MARL
Table 2 provides a detailed side-by-side comparison of our approach with recent sampling-based methods for networked MARL [22, 26]. The comparison highlights key differences in sampling targets, distribution types, locus of adaptivity, and first-hop inclusion probabilities.
Table 2: Comparison of sampling strategies in networked MARL. Our method learns context-dependent, per-edge distributions optimized jointly with the policy, in contrast to predefined or uniform sampling approaches.
Aspect Lin et al. [22] Anand & Qu [26] BayesG (Ours)
Distribution for sampling Exogenous, fixed distribution $\mathcal{D}$ over active link sets $(L_{t}^{s},L_{t}^{r})$, sampled i.i.d.; assumes distance truncation/decay Fixed, uniform distribution over size-$k$ subsets of local agents for value estimation and action Learned, per-agent variational posterior $q(Z_{i};\phi_{i})$ over 1-hop edges; state/trajectory-dependent, trained end-to-end (ELBO)
Sampling target Global edges that are "active" for transitions ($L^{s}$) and rewards ($L^{r}$) at each step Agents (a subset of locals) for the single global agent to aggregate over (not edges) Binary edge mask on each agent’s 1-hop ego-graph (which neighbors to "listen to" for policy/communication)
Locus of dynamic topology Dynamics-level stochasticity: new active link sets each step; multi-hop influence via chains across time Agent-level subsampling: the global agent’s view changes as different locals are sampled; topology effectively star-like, not edge-level Policy/representation-level adaptivity: each agent samples a task-adaptive 1-hop subgraph each step; no multi-hop rewiring (multi-hop only via repeated local passes)
1-hop inclusion probability Inclusion of $(j\!\to\!i)$ is the marginal of $\mathcal{D}$ (fixed/model-assumed; not learned or state-adaptive) No edge-level 1-hop probability; agent-level inclusion is $k/n$ (uniform, fixed) per step for the global agent For neighbor $j$ of agent $i$, inclusion $\Pr[z_{ij}{=}1]=\sigma(\phi_{ij}(\cdot))$ is learned and context-dependent (from $q$)
These distinctions highlight that while all three methods employ sampling for scalability, they differ fundamentally in whether the sampling distribution is learned, what structural elements are sampled, and where adaptivity occurs in the system architecture.
Appendix I Limitations
@@
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.